According to the management entrenchment theory, managers choose capital structure so as to preserve their control of the firm. On the one hand, debt is costly for managers because risk losing control in the event of default. On the other hand if they do not take advantage of the tax shield provided by debt, they risk losing control through a hostile takeover. Suppose a firm expects to generate free cash flows of 86 million per year, and the discount rate for these cash flows is 9%. The firm pays a tax rate of 40%. A raider is poised to take over the firm and finance it with 585 million in permanent debt. The raider will generate the same cash flows and the takeover attempt will be successful if the raider can offer a premium of 21% over the current value of the firm. According to the managerial entrenchment hypothesis, what level of permanent debt will the firm choose?
Answer #1

Unlevered value= CF/Discount rate

= 86m/9%

= 955.56 m

Levered value with the raider = Unlevered value+ Tax*Debt

= 955.56+ 40%*585

=1189.56

Minimum levered value required to prevent the raid = Levered value/ (1+Premium)

= 1189.56/ (1+0.21)

=983.11

Minimum tax shield = 983.11- 955.56 = 27.55

Level of debt required = 27.55/40% = 68.875 million